The city of Charleston had the following sales of water for the selected months of 2017:

Month Sales

February 60,000

March 45,000

April 65,000

May 43,350

June 110,000

July 120,000

All sales are on credit. Historically, 60 percent is collected in the month of sale, 30 percent during the first month following the sale, and 10 percent in the second month following the sale. Water purchases by month are as follows:

Month Sales

February 45,000

March 45,000

April 40,000

May 59,750

June 52,500

July 90,000

Water is purchased in the month of sale. All purchases are paid during the month following the purchase. Operating costs are $18,000 and everything is paid in cash except for depreciation, which totals $8,000 a month. The city plans on purchasing some new equipment in May for $25,000 in exchange for a note payable. The April 1 cash balance is expected to be $5,000. The city must maintain a minimum cash balance of $10,500, and money can be borrowed from a local bank in increments of $1,000. The city borrows money at the beginning on the first day of the month and repays loans and interest on the last day of the month. The bank charges the city an annual interest rate of 15%.

Required: Prepare a cash budget for April, May, June and in for the quarter, and based on your answer complete the following table: Round to the nearest dollar and DO NOT enter decimals, or commas and if a zero needs to be entered, enter "0".

April May June Quarter
Beginning Cash Balance 5,000 10,500 10,510 5,000
Cash Collections
Water Disbursements
Operating costs
Borrowings
Repayments
Interest
Ending Cash Balance 10,500 10,510 24,215 24,215
Answer #1
The city of Charleston
Calculation
Cash Budget April May June Quarter April May June
Beginning Cash balance            5,000          10,500          10,510            5,000
Cash Collections          58,500          50,010          85,505        194,015 Workings below
Water Disbursements        (45,000)        (40,000)        (59,750)     (144,750)
Operating costs        (10,000)        (10,000)        (10,000)        (30,000) =18,000-8,000 =18,000-8,000 =18,000-8,000
Borrowings            2,000                   -                     -              2,000 =10,500-8,500
Repayments                   -                     -            (2,000)          (2,000)
Interest                   -                     -                  (50)                (50) =2,000*15%*2/12
Ending Cash balance          10,500          10,510          24,215          24,215
Workings
Cash Receipts Budget April May June Quarter April May June
Sales          65,000          43,350        110,000        218,350
Cash Receipts from:
February sales            6,000                   -                     -              6,000 =60,000*10%
March sales          13,500            4,500                   -            18,000 =45,000*30% =45,000*10%
April sales          39,000          19,500            6,500          65,000 =65,000*60% =65,000*30% =65,000*10%
May sales                   -            26,010          13,005          39,015 =43,350*60% =43,350*30%
June sales                   -                     -            66,000          66,000 =110,000*60%
Total Cash Receipts          58,500          50,010          85,505        194,015